How would I write the following so the coefficient is in front?
1. (3y)8
2. 3(4b)

Found 3 solutions by Boreal, MowMow, ikleyn:
Answer by Boreal(15235)

Multiply the coefficients together and put them first
24y
The other is 12b

Answer by MowMow(42)

1. (3y)8 = (3^8)y^8 = 6561y^8

2. 3(4b) = 12b

Answer by ikleyn(53570)

.
How would I write the following so the coefficient is in front?
1. (3y)8
2. 3(4b)
~~~~~~~~~~~~~~~~~~~~~~~~

The answer (1) in the post by @MowMow is INCORRECT.

The correct answer to (1) is (3*8)*y = 24y.